# Blending Data-Driven Priors in Dynamic Games

### v2

@@ Abstract
As intelligent robots like autonomous vehicles become increasingly deployed in the presence of people, the extent to which these systems should leverage model-based game-theoretic planners versus data-driven policies for safe, interaction-aware motion planning remains an open question. Existing dynamic game formulations assume all agents are task-driven and behave optimally. However, in reality, humans tend to deviate from the decisions prescribed by these models, and their behavior is better approximated under a noisy-rational paradigm. In this work, we investigate a principled methodology to blend a data-driven reference policy with an optimization-based game-theoretic policy. We formulate KLGame, a type of non-cooperative dynamic game with Kullback-Leibler (KL) regularization with respect to a general, stochastic, and possibly multi-modal reference policy. Our method incorporates, for each decision maker, a tunable parameter that permits modulation between task-driven and data-driven behaviors. We propose an efficient algorithm for computing multi-modal approximate feedback Nash equilibrium strategies of KLGame in real time. Through a series of simulated and real-world autonomous driving scenarios, we demonstrate that KLGame policies can more effectively incorporate guidance from the reference policy and account for noisily-rational human behaviors versus non-regularized baselines.
@@ I Introduction
Two diverging approaches have recently emerged for planning in the presence of nearby agents. Dynamic game theory [1], which models interactions as a multi-agent optimal control problem, allows computing a task-optimal plan simultaneously for all agents in the scene. However, while these methods can effectively capture high-level rational behaviors, other agents do not always act rationally—i.e., they do not necessarily pick the game-theoretically optimal action. Moreover, gradient-based game-theoretic solvers may struggle to model multi-modal, mixed-strategy (stochastic) interactions, wherein distinct high-level behaviors may arise from the same objectives, leading to several local optima in the space of rewards. Conversely, behavior cloning methods provide a framework for end-to-end learning of multi-modal trajectory distributions from data, but may yield suboptimal or unsafe predictions.
Figure 1: Example trajectories for a game-theoretic and a data-driven motion planner (top row) versus the trajectories generated by KLGame (bottom row), where the ego vehicle interacts with a distracted pedestrian crossing the road. Different trajectories are shown for the KLGame to illustrate the influence of the regularization weight, $\lambda$, on the policy blending. $\lambda=0$ does not introduce any policy alignment and purely follows a game-theoretic planning solution. As $\lambda$ increases, the trajectory becomes aligned with the policy dictated by the data priors.
@@ II-A Game-Theoretic Motion Planning
Game-theoretic motion planning is a popular choice for modeling multi-agent non-cooperative interactions, such as autonomous driving [5, 4, 7, 8, 9], crowd navigation [10], distributed systems [11, 12], drone racing [13], and shared control [14]. A central focus of game-theoretic planning are local Nash equilibria (LNE) [1], in which no agent is unilaterally incentivized to deviate from their strategy. LNEs have been studied extensively in autonomous driving as a model for interaction in merging [15, 16, 17], highway overtaking [18, 19, 20], and racing [21, 8]. However, in many real-world settings, agents are not perfectly rational and may deviate significantly from actions they believe are optimal due to distractions [22], imperfect information [23, 24], or poor human-machine interfacing [25]. In this work, we adopt the principle of bounded rationality [26, 27, 28, 29], in which agents act rationally with likelihood proportional to the distribution of utilities over actions.
@@
While LNEs provide a mechanism for predicting and analyzing multi-agent interactions, multiple LNEs may arise depending on the initial condition of the joint (multi-agent) system [1]. Scenarios where multiple equilibria exist are called multi-modal due to the existence of distinct (and possibly stochastic) outcomes, the selection of which might depend on individual preferences. Multi-modality in planning has recently become a popular object of study for modeling high-level decision making in games. Peters et al. [30] introduce an inference framework for maximum a posteriori equilibria aligned control by randomly selecting seed strategies and solving multiple seeds in parallel to an equilibrium, to then normalize their weights in a manner similar to a particle filter. So et al. [31, 32] show that equilibria can be efficiently clustered and inferred using a Bayesian update, which can then be used to hedge different strategies using a QMDP-style [33] cost estimate. Peters et al. [34] show that multiple game-theoretic contingencies can be solved, with an arbitrary branching time, using mixed-complementarity programming. Recent work [9] uses implicit dual control to tractably compute an active information gathering policy for a class of partially-observable stochastic games, while preserving the multi-modal information encoded in the robot’s belief states with scenario optimization. Follow-up work [24] extends this idea to a high-dimensional adversarial setting using deep reinforcement learning to synthesize the robot’s policy at scale.
@@
Multi-modality in planning can arise from uncertain or unknown objectives. For example, in a merge setting, which player prefers to go first may directly alter the outcome of interaction, and this objective uncertainty may lead to indecision. Cleac’h et al. [18] introduce a planning for uncertain linear cost models by estimating cost parameters using an unscented Kalman filter. Chen et al. [35] introduce an MPC-based framework for interactive trajectory optimization using deep-learning predictions as input. Hu et al. [36] study emergent coordination in multi-agent interactions when cost parameters follow an opinion-dynamic interaction over a graph. Recent work [37] uses differentiable optimization to learn open-loop mixed strategies for non-cooperative games based on trajectory data. Liu et al. [17] extend this idea to compute generalized equilibria, a setting where hard constraints are present, and use deep learning to bolster the computation performance of the game solver. Another follow-up work [38] further studies the setting of learning a feedback Nash equilibrium strategy. In the potential game setting, Diehl et al. [39] show that multiple open-loop Nash equilibria can be computed in parallel using deep-learned game parameters.
@@ II-B Stochastic Optimal Control and Dynamic Games
In contrast to greedy-optimal, deterministic behavior predicted by some classical control-theoretic planners, stochastic behavior has been exhibited in many real-world settings, such as biological muscle control [40], animal foraging [41], and urban driving [42]. Stochastic optimal control (SOC) provides a mechanism for solving and explaining stochastic optimal policies through accommodating uncertainty in decision making. Energy-based methods [43, 44, 45, 46] are a well-studied technique for solving optimal policies by analyzing utility functions similar to the total energy in statistical physics. KL control [47, 48, 49, 50] is a branch of energy-based methods that studies solutions for SOC problems where the control cost is augmented with a Kullback-Leibler divergence regularizer term that tries to keep the resultant policy close to some reference policy, which may have different uses depending on the task. For example, the reference policy could minimize control effort [47], guide exploration [51], or incorporate human feedback [52].
@@ II-C Multi-Modal Multi-Agent Motion Prediction
In sample-based prediction, maintaining the diversity of samples to represent distinct outcomes is an active area of research. Recent works [73, 74] emphasize multi-modality by incorporating rollouts and fictitious play to increase diversity in the prediction framework. Farthest Point Sampling (FPS) [75, 76], Non-Maximum Suppression (NMS) [77], and/or neural adaptive sampling [78] provide methods for encouraging outcome diversity using pairwise distances between trajectories as a metric. Sample diversity makes prediction methods a popular data-driven choice for human behavior. Prediction models have been incorporated in planning through an integrated prediction and planning approach, typically a model-predictive controller that seeks to avoid other agents [3, 79]. In [2], joint motion forecasts are incorporated as a constraint in the motion planning framework, allowing the planning agent to better understand high-dimensional scene data (such as the map) through low-dimensional action selection. In [80], a parameter-shared cost model is learned from data for all agents in the scene, allowing for scene-dependent cost formulation.
@@ III Problem Formulation
where $x_{t}\in\mathcal{X}$ is the joint state of the dynamical system and $u_{t}:=(u_{t}^{1},...,u_{t}^{N})\in\prod_{i\in[N]}\mathcal{U}^{i}$ is the joint control input. Each player $i\in[N]$ applies an action according to a stochastic policy ${\pi}_{t}^{i}({u}^{i}_{t}|{x}_{t})$ at each time step $t$. For notational convenience, we occasionally will write ${\pi}_{t}^{i}({u}^{i}_{t}|{x}_{t})={\pi}_{t}^{i}(\cdot|{x}_{t})$.
@@
For the KL divergence to be well-defined, we require that ${\pi}_{t}^{i}(\cdot|x_{t})$ is absolutely continuous with respect to $\tilde{\pi}_{t}^{i}(\cdot|x_{t})$, i.e., ${\pi}^{i}_{t}<<\tilde{\pi}_{t}^{i}$ The scalar ${\lambda}^{i}\geq 0$ provides a natural “tuning knob” for the stochastic dynamic game, balancing the preference between optimizing the player’s nominal performance objectives, encoded in $l^{i}$, and aligning its policy with the reference policy, ${\tilde{\pi}}^{i}$. We refer to this new class of dynamic game as KLGame, which can be compactly formulated as:
@@
Herein, we use $\neg i$ to denote all agents except for $i$, and the superscript $(\cdot)^{*}$, as an indicator of optimality. We now introduce the main technical approach for seeking FBNE of KLGame.
@@ IV-A Dynamic Programming Formulation
where ${V}^{i}({x})=\inf_{\pi^{i}}J^{i}(\pi^{i})$ denotes the optimal value function and ${\mathcal{Q}}^{i}({x},{u}):=l^{i}({x},{u})+V^{i\prime}(f({x},{u}))$ is the Q-value function. In the reminder of this section, we show that (5) admits an optimal solution of a particular structure.
@@ IV-B Global FBNE: The Linear-Quadratic Gaussian Case
Even without the KL regularizer in (2), seeking a local feedback Nash equilibrium of a non-LQ trajectory game is generally intractable [84]. However, if the system dynamics are linear, the cost functions are convex quadratic, and the reference policy is Gaussian, the global Nash equilibrium can be computed efficiently. We call KLGame in this special case the KL-regularized Linear-Quadratic Gaussian (KL-LQG) game, whose equilibrium properties are analyzed in what follows.
@@
    The dynamics follow the functional form
    $x_{t+1}=A_{t}x_{t}+\sum_{i\in[N]}B_{t}^{i}u^{i}_{t}+{d}_{t},$
    where $x_{0}\sim{\mathcal{N}}(\mu_{x_{0}},\Sigma_{x_{0}})$, ${d}_{t}\sim{\mathcal{N}}(0,\Sigma_{{d}})$,
@@
    The costs have the functional form
@@
    The reference policies are Gaussian, i.e., ${\tilde{\pi}}^{i}_{t}\sim{\mathcal{N}}(\tilde{\mu}^{i}_{t},\tilde{\Sigma}^{i}_{t})$ for all $t\in[T]$ and $i\in[N]$.
@@
$\Sigma^{i*}_{t}$ $=\left[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\frac{1}{{\lambda}^{i}}}\left(R^{ii}_{t}+{B^{i}_{t}}^{T}Z^{i}_{t+1}B^{i}_{t}\right){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}+\left(\tilde{\Sigma}^{i}_{t}\right)^{-1}}\right]^{-1},$
@@
$=B_{t}^{i\top}z^{i}_{t+1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}-{\lambda}^{i}(\tilde{\Sigma}^{i}_{t})^{-1}\tilde{\mu}^{i}_{t}}.$
@@
$Z_{t}^{i}=$ $Q_{t}^{i}+\sum_{j\in[N]}(K_{t}^{j})^{T}R_{t}^{ij}K_{t}^{j}+F_{t}^{T}Z_{t+1}^{i}F_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}+{\lambda}^{i}K^{i\top}_{t}(\tilde{\Sigma}^{i}_{t})^{-1}K^{i}_{t}},$ (10)
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}+{\lambda}^{i}K^{i\top}_{t}(\tilde{\Sigma}^{i}_{t})^{-1}(\kappa^{i}_{t}-\tilde{\mu}^{i}_{t})},$
@@
  • The global FBNE must be a mixed strategy, taking the form of a unimodal Gaussian.
@@
Recall from [4] that an ECE in the LQ case is also a unimodal Gaussian. When $\min~{}\sigma(\tilde{\Sigma}^{i}_{t})\rightarrow\infty$, the contribution from the reference policy (red terms) in the covariance $\Sigma^{i*}_{t}$ and Riccati equation vanishes, and the resulting covariance and Riccati equation match that of the maximum-entropy dynamic game. ∎
@@
$=B^{i\top}Z^{i}_{t+1}A{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}+{\lambda}^{i}(\tilde{\Sigma}^{i}_{t})^{-1}\tilde{K}^{i}_{t}},$
@@
$=B^{i\top}z^{i}_{t+1}{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}+{\lambda}^{i}(\tilde{\Sigma}^{i}_{t})^{-1}\tilde{\kappa}^{i}_{t}}.$
@@
${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}+{\lambda}^{i}(K^{i}_{t}+\tilde{K}^{i}_{t})^{\top}(\tilde{\Sigma}^{i}_{t})^{-1}(K^{i}_{t}+\tilde{K}^{i}_{t})},$
@@
${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}+{\lambda}^{i}(K^{i}_{t}+\tilde{K}^{i}_{t})^{\top}(\tilde{\Sigma}^{i}_{t})^{-1}(\kappa^{i}_{t}+\tilde{\kappa}^{i}_{t})},$
@@
IV-C Solving General KLGame Planning Porblems

### v3

@@ Abstract
As intelligent robots like autonomous vehicles become increasingly deployed in the presence of people, the extent to which these systems should leverage model-based game-theoretic planners versus data-driven policies for safe, interaction-aware motion planning remains an open question. Existing dynamic game formulations assume all agents are task-driven and behave optimally. However, in reality, humans tend to deviate from the decisions prescribed by these models, and their behavior is better approximated under a noisy-rational paradigm. In this work, we investigate a principled methodology to blend a data-driven reference policy with an optimization-based game-theoretic policy. We formulate KLGame, an algorithm for solving non-cooperative dynamic game with Kullback-Leibler (KL) regularization with respect to a general, stochastic, and possibly multi-modal reference policy. Our method incorporates, for each decision maker, a tunable parameter that permits modulation between task-driven and data-driven behaviors. We propose an efficient algorithm for computing multi-modal approximate feedback Nash equilibrium strategies of KLGame in real time. Through a series of simulated and real-world autonomous driving scenarios, we demonstrate that KLGame policies can more effectively incorporate guidance from the reference policy and account for noisily-rational human behaviors versus non-regularized baselines. Website with additional information, videos, and code: https://kl-games.github.io/.
@@ I Introduction
Figure 1: Our proposed approach can seamlessly integrate data-driven policies with optimization-based dynamic game solutions. Top: Data-driven behavior predictors, such as transformer-based methods, provide marginal (i.e., single-agent) priors for human motion prediction, but may struggle to model closed-loop, multi-agent interactions. Middle: Our proposed method, KLGame, allows a robot to incorporate guidance from data-driven rollouts while performing online game-theoretic planning in closed-loop. Bottom: KLGame uses a tunable parameter $\lambda$ that modulates behaviors on a spectrum: $\lambda=0$ gives a deterministic dynamic game (task-optimal), and $\lambda\rightarrow\infty$ gives multi-modal behavior cloning. We call this tunability policy blending.
@@
Two diverging approaches have recently emerged for planning in the presence of nearby agents. Dynamic game theory [1], which models interactions as a multi-agent optimal control problem, allows computing a task-optimal plan simultaneously for all agents in the scene. However, while these methods can effectively capture high-level rational behaviors, other agents do not always act rationally—i.e., they do not necessarily pick the game-theoretically optimal action. Moreover, gradient-based game-theoretic solvers may struggle to model multi-modal, mixed-strategy (stochastic) interactions, wherein distinct high-level behaviors may arise from the same objectives, leading to several local optima in the space of rewards. Conversely, behavior cloning methods provide a framework for end-to-end learning of multi-modal trajectory distributions from data, but may yield suboptimal or unsafe predictions. Such data-driven methods exploit recent advances in autonomous vehicle perception systems by synthesizing high-dimensional sensor inputs, including a road map, traffic light states, and LIDAR data. While imitation learning can provide a flexible and efficient approximation for interactions modeled by dynamic game theory, predictions may be suboptimal or unsafe.
@@ II-A Game-Theoretic Motion Planning
Game-theoretic motion planning is a popular choice for modeling multi-agent non-cooperative interactions, such as autonomous driving [5, 4, 7, 8, 9], crowd navigation [10], distributed systems [11, 12], drone racing [13], and shared control [14]. A central focus of game-theoretic planning is local Nash equilibria (LNE) [1], in which no agent is unilaterally incentivized to deviate from their strategy. LNEs have been studied extensively in autonomous driving as a model for interaction in merging [15, 16, 17], highway overtaking [18, 19, 20], and racing [21, 8]. However, in many real-world settings, agents are not perfectly rational and may deviate significantly from actions they believe are optimal due to distractions [22], imperfect information [23, 24], or poor human-machine interfacing [25]. In this work, we adopt the principle of bounded rationality [26, 27, 28, 29], in which agents act rationally with likelihood proportional to the distribution of utilities over actions.
@@
While LNEs provide a mechanism for predicting and analyzing multi-agent interactions, multiple LNEs may arise depending on the initial condition of the joint (multi-agent) system [1]. Scenarios where multiple equilibria exist are called multi-modal due to the existence of distinct (and possibly stochastic) outcomes, the selection of which might depend on individual preferences. Multi-modality in planning has recently become a popular object of study for modeling high-level decision-making in games. Peters et al. [30] introduce an inference framework for maximum a posteriori equilibria aligned control by randomly selecting seed strategies and solving multiple seeds in parallel to an equilibrium, to then normalize their weights in a manner similar to a particle filter. So et al. [31, 32] show that equilibria can be efficiently clustered and inferred using a Bayesian update, which can then be used to hedge different strategies using a QMDP-style [33] cost estimate. Peters et al. [34] show that multiple game-theoretic contingencies can be solved, with an arbitrary branching time, using mixed-complementarity programming. Recent work [9] uses implicit dual control to tractably compute an active information gathering policy for a class of partially-observable stochastic games, while preserving the multi-modal information encoded in the robot’s belief states with scenario optimization. Follow-up work [24] extends this idea to a high-dimensional adversarial setting using deep reinforcement learning to synthesize the robot’s policy at scale.
@@
Multi-modality in planning can arise from uncertain or unknown objectives. For example, in a merging setting, which player prefers to go first may directly alter the outcome of the interaction, and this objective uncertainty may lead to indecision. Cleac’h et al. [18] introduce a motion planning method for uncertain linear cost models by estimating cost parameters using an unscented Kalman filter. Chen et al. [35] introduce an MPC-based framework for interactive trajectory optimization using deep-learning predictions as input. Hu et al. [36] study emergent coordination in multi-agent interactions when cost parameters follow an opinion-dynamic interaction over a graph. Recent work [37] uses differentiable optimization to learn open-loop mixed strategies for non-cooperative games based on trajectory data. Liu et al. [17] extend this idea to compute generalized equilibria, a setting where hard constraints are present, and use deep learning to bolster the computation performance of the game solver. Another follow-up work [38] further studies the setting of learning a feedback Nash equilibrium strategy. In the potential game setting, Diehl et al. [39] show that multiple open-loop Nash equilibria can be computed in parallel using deep-learned game parameters.
@@ II-B Stochastic Optimal Control and Dynamic Games
In contrast to greedy-optimal, deterministic behavior predicted by some classical control-theoretic planners, stochastic behavior has been exhibited in many real-world settings, such as biological muscle control [40], animal foraging [41], and urban driving [42]. Stochastic optimal control (SOC) provides a mechanism for solving and explaining stochastic optimal policies through accommodating uncertainty in decision-making. Energy-based methods [43, 44, 45, 46] are a well-studied technique for solving optimal policies by analyzing utility functions similar to the total energy in statistical physics. KL control [47, 48, 49, 50] is a branch of energy-based methods that studies solutions for SOC problems where the control cost is augmented with a Kullback-Leibler divergence regularizer term that tries to keep the resultant policy close to some reference policy, which may have different uses depending on the task. For example, the reference policy could minimize control effort [47], guide exploration [51], or incorporate human feedback [52].
@@ II-C Multi-Modal Multi-Agent Motion Prediction
In sample-based prediction, maintaining the diversity of samples to represent distinct outcomes is an active area of research. Recent works [73, 74] emphasize multi-modality by incorporating rollouts and fictitious play to increase diversity in the prediction framework. Farthest Point Sampling (FPS) [75, 76], Non-Maximum Suppression (NMS) [77], and/or neural adaptive sampling [78] provide methods for encouraging outcome diversity using pairwise distances between trajectories as a metric. Sample diversity makes prediction methods a popular data-driven choice for human behavior.
Prediction models have been incorporated in planning through an integrated prediction and planning approach, typically a model-predictive controller that seeks to avoid other agents [3, 79]. In [2], joint motion forecasts are incorporated as a constraint in the motion planning framework, allowing the planning agent to better understand high-dimensional scene data (such as the map) through low-dimensional action selection. In [80], a parameter-shared cost model is learned from data for all agents in the scene, allowing for scene-dependent cost formulation.
Generative modeling of traffic scenarios is also capable of compositing learned behavior behaviors with model-based objectives. Recent works [81, 82] utilize signal temporal logic to guide the diffusion model generating desired behaviors. VBD [83] applies a game-theoretic gradient descent-ascent guidance strategy to generate interactive safety-critical scenarios with a diffusion model.
In this work, we provide a principled approach to integrate the multi-modality of motion predictors with closed-loop game-theoretic motion planning. Our framework complements end-to-end learned methods by allowing an agent to consider multiple motion modes and compute an optimal policy for each one, rather than needing to select a mode a-priori. Furthermore, by specifying additional incentives such as safety, KLGame allows a robot to reason strategically over time in a closed-loop manner, making it more robust to unmodelled scenarios through feedback when compared to open-loop end-to-end predictions.
@@ III Problem Formulation
Notation. We will take $T$ as the planning horizon and define $[T]:=\{1,...,T\}$. Similarly, if there are $N$ players we will use $[N]:=\{1,...,N\}$ to refer to the set containing all the players’ indices. Throughout the paper we will use subscripts to refer to time and superscripts for player. For example, $u_{t}^{i}$ will stand for the action of player $i$ at time $t$. For a matrix $M$ we use $M\succ 0$ (or $M\succeq 0$) to indicate M is positive (or semi-positive) definite. We use $\sigma(M)$ to indicate the set of singular values of $M$.
@@
where $x_{t}\in\mathcal{X}$ is the joint state of the dynamical system and $u_{t}:=(u_{t}^{1},...,u_{t}^{N})$ is the joint control input. Each player $i\in[N]$ applies an action according to a stochastic policy ${\pi}_{t}^{i}({u}^{i}_{t}|{x}_{t})$ at each time step $t$.
@@
For the KL divergence to be well-defined, we require that ${\pi}_{t}^{i}(\cdot|x_{t})$ is absolutely continuous with respect to $\tilde{\pi}_{t}^{i}(\cdot|x_{t})$, i.e., $\tilde{\pi}_{t}^{i}(\cdot|x_{t})=0\implies{\pi}_{t}^{i}(\cdot|x_{t})=0$ [84]. We denote absolute continuity as ${\pi}^{i}_{t}<<\tilde{\pi}_{t}^{i}$.
The scalar ${\lambda}^{i}\geq 0$ provides a natural “tuning knob” for the stochastic dynamic game, balancing the preference between optimizing the player’s nominal performance objectives, encoded in $l^{i}$, and aligning its policy with the reference policy, ${\tilde{\pi}}^{i}$. We refer to this new class of dynamic game as KLGame, which can be compactly formulated as:
@@
We now introduce the main technical approach for seeking FBNE of KLGame.
@@ IV-A Dynamic Programming Formulation
where ${V}^{i}({x})=\inf_{\pi^{i}}J^{i}(\pi^{i})$ denotes the optimal value function and ${\mathcal{Q}}^{i}({x},{u}):=l^{i}({x},{u})+V^{i\prime}(f({x},{u}))$ is the Q-value function. In the remainder of this section, we show that (5) admits an optimal solution of a particular structure.
@@ IV-B Global FBNE: The Linear-Quadratic Gaussian Case
Even without the KL regularizer in (2), seeking a local feedback Nash equilibrium of a non-LQ trajectory game is generally intractable [88]. However, if the system dynamics are linear, the cost functions are convex quadratic functions of the state (LQ), and the reference policy is Gaussian, the global Nash equilibrium can be computed efficiently. Specifically, the system dynamics are linear when
$x_{t+1}=A_{t}x_{t}+\sum_{i\in[N]}B_{t}^{i}u^{i}_{t}+{d}_{t},$ (8)
where $A_{t},B_{t}^{i}$ are the state-transition matrices and ${d}_{t}\sim{\mathcal{N}}(0,\Sigma_{{d}})$ Gaussian noise with zero mean and covariance $\Sigma_{d}$. The stage cost of each player is a quadratic function of state and control, i.e., $l^{i}(x_{t},u_{t})=x_{t}^{\top}Q^{i}_{t}x_{t}+\sum_{j\in[N]}{u_{t}^{j}}^{\top}R_{t}^{ij}u_{t}^{j}$, where $Q^{i}$ and $R^{ij}$ are positive semidefinite matrices of appropriate dimension penalizing state and control authority. The reference policies ${\tilde{\pi}}^{i}$ are Gaussian if
${\tilde{\pi}}^{i}_{t}={\mathcal{N}}(\tilde{\mu}^{i}_{t},\tilde{\Sigma}^{i}_{t})$ (9)
for all $t\in[T]$ and $i\in[N]$.
We call KLGame, in this special case, the KL-regularized Linear-Quadratic Gaussian (KL-LQG) game. Our goal is to find, for all agents $i$, the global FBNE strategy $\pi^{i*}$ of a KL-LQG game. We show in the following theorem that $\pi^{i*}$ is Gaussian with a state-dependent mean $\mu^{i*}=\mu^{i*}(x)$, which depends linearly on $x$, and a state-independent covariance $\Sigma^{i*}.$ Furthermore, we show that $\mu^{i*}$ and $\Sigma^{i*}$ are parameterized by the dynamics ($A$ and $B$), task costs ($Q$ and $R$), reference policy ($\tilde{\mu}$ and $\tilde{\Sigma}$), and value function parameters.
@@
    The dynamics follow the functional form of (8), where $x_{0}\sim{\mathcal{N}}(\mu_{x_{0}},\Sigma_{x_{0}})$, ${d}_{t}\sim{\mathcal{N}}(0,\Sigma_{{d}})$,
@@
    The costs, introduced in (2), have the functional form
@@
    The reference policies are Gaussian, as in (9).
@@
$\Sigma^{i*}_{t}$ $=\left[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{1}{{\lambda}^{i}}}\left(R^{ii}_{t}+{B^{i}_{t}}^{T}Z^{i}_{t+1}B^{i}_{t}\right){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+\left(\tilde{\Sigma}^{i}_{t}\right)^{-1}}\right]^{-1},$
where $(K^{i}_{t},\kappa^{i}_{t})$, the state-feedback matrix and constant, are given by solving the coupled KL-regularized Riccati equation:
@@
$=B_{t}^{i\top}z^{i}_{t+1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-{\lambda}^{i}(\tilde{\Sigma}^{i}_{t})^{-1}\tilde{\mu}^{i}_{t}}.$
@@
$Z_{t}^{i}=$ $Q_{t}^{i}+\sum_{j\in[N]}(K_{t}^{j})^{T}R_{t}^{ij}K_{t}^{j}+F_{t}^{T}Z_{t+1}^{i}F_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+{\lambda}^{i}K^{i\top}_{t}(\tilde{\Sigma}^{i}_{t})^{-1}K^{i}_{t}},$ (12)
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+{\lambda}^{i}K^{i\top}_{t}(\tilde{\Sigma}^{i}_{t})^{-1}(\kappa^{i}_{t}-\tilde{\mu}^{i}_{t})},$
@@
  • The global FBNE must be a mixed strategy, taking the form of an unimodal Gaussian.
@@
Recall from [4] that an ECE in the LQ case is also an unimodal Gaussian. When $\min~{}\sigma(\tilde{\Sigma}^{i}_{t})\rightarrow\infty$, the contribution from the reference policy (red terms) in the covariance $\Sigma^{i*}_{t}$ and Riccati equation vanishes, and the resulting covariance and Riccati equation match that of the maximum-entropy dynamic game. ∎
@@
$=B^{i\top}Z^{i}_{t+1}A{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+{\lambda}^{i}(\tilde{\Sigma}^{i}_{t})^{-1}\tilde{K}^{i}_{t}},$
@@
$=B^{i\top}z^{i}_{t+1}{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+{\lambda}^{i}(\tilde{\Sigma}^{i}_{t})^{-1}\tilde{\kappa}^{i}_{t}}.$
@@
${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+{\lambda}^{i}(K^{i}_{t}+\tilde{K}^{i}_{t})^{\top}(\tilde{\Sigma}^{i}_{t})^{-1}(K^{i}_{t}+\tilde{K}^{i}_{t})},$
@@
${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}+{\lambda}^{i}(K^{i}_{t}+\tilde{K}^{i}_{t})^{\top}(\tilde{\Sigma}^{i}_{t})^{-1}(\kappa^{i}_{t}+\tilde{\kappa}^{i}_{t})},$
@@
IV-C Solving General KLGame Planning Problems